How old is Alex if his age six years from now with be twice his age 7 years ago
Answer by Cromlix(4381)

Hi there,
How old is Alex if his age six years from
now will be twice his age 7 years ago?
Make Alex's age = x
His age six years from now will be
x + 6
twice his age 7 years ago
2(x - 7)
Make two equations equal one another.
x + 6 = 2(x - 7)
x + 6 = 2x - 14
collect like terms
x - 2x = - 14 - 6
-x = -20
x = 20
Alex is 20 years old.
Hope this helps :-)